OpenStudy (anonymous):

PLEASE SHOW ALL WORK I'M SO LOST AND NEED TO UNDERSTAND THIS! 1. Part 1: Solve the following equation using one of the methods listed below. Show all work. (4 points) Part 2: Using complete sentences, explain why you chose the method you used. (2 points) Equation: 0 = x2 + 6x + 3 Methods: Completing the Square Factoring Quadratic Formula Graphing 2. Jon’s new mini computer monitor has a 20 cm by 24 cm LCD screen with a plastic trim border. The total area of the monitor and border is 672 cm2. How wide is the trim around the monitor?

OpenStudy (anonymous):

\(0 = x^2 + 6x + 3 \)does not factor using integers, but \(6\) is even so completing the square easiest

OpenStudy (anonymous):

\[ x^2 + 6x + 3=0 \]subtract 3\[x^2+6x=-3\]half of 6 is 3, and 3 squared is 9 so write \[(x+3)^2=-3+9=6\]take the square root, don''t forget the \(\pm\) write \[x+3=\pm\sqrt{6}\]solve for \(x\) get \[x=-3\pm\sqrt{6}\]

OpenStudy (anonymous):

What about number 1?